### v2

@@ Abstract
The research findings highlight the effectiveness of the proposed approach, with the CANINE demonstrating superior performance in analyzing page titles and the RoBERTa excelling in evaluating page content. The fusion of two NLP and one MLP model, termed MultiText-LP, achieves impressive results, yielding a 96.80 F1 score and a 97.18 accuracy score on our research dataset. Furthermore, our approach outperforms existing methods on the CatchPhish HTML dataset, showcasing its efficacies.
@@ 1 Introduction
The Internet Revolution began in the late 20th century and has significantly changed how we communicate, work, and live [pencarelli2020digital]. It has connected people worldwide, providing unprecedented access to information, resources, and opportunities. However, along with the benefits, there is a growing challenge of internet dependence and an increase in cyber-attacks [kamruzzaman2016plight].
One notable threat in this digital age is phishing, a deceitful practice where cybercriminals trick individuals into revealing sensitive information like usernames, passwords, or financial details [katkuri2018indian]. Phishing techniques have become sophisticated, utilizing deceptive emails, fake websites, and malicious software. Staying vigilant and informed is crucial as cybercriminals continuously evolve their strategies.
Phishing has become the most common form of cybercrime today. According to the Anti-Phishing Working Group (APWG), in the second quarter of 2023, 1,286,208 phishing attacks were observed—the third-highest quarterly total on record [apwg2023]. The average wire transfer amount in Business Email Compromise (BEC) attacks surged to \$293,359, a 57% increase from Q1. Additionally, an estimated 3.4 billion phishing emails are sent daily, with 1,350,037 reported attacks in Q4 2022 alone. Notably, 95% of social engineering attack motivations are financially driven, emphasizing the high stakes involved [apwg2023].
Various software-based detection techniques are utilized in the ongoing fight against phishing attacks. Blacklists [prakash2010phishnet] and whitelists [jain2016novel] serve as standard tools, but their effectiveness may be limited in identifying new or subtly modified phishing sites. Machine learning-based approaches leverage heuristic features like URL [rao2020catchphish] [prabakaran2023enhanced], HTML content [jain2019machine] [rao2022wordembedding] [opara2020htmlphish], website traffic, search engine data, and WHOIS records for classification, enhancing overall efficiency. Third-party services contribute valuable data to the detection process, including page rank, search engine indexing, and WHOIS [cheng2022detecting] [kuyama2016method]. Visual similarity-based heuristics compare new websites with pre-stored signatures based on screenshots, font styles, images, page layouts, and logos [abdelnabi2020visualphishnet] [chen2020intelligent].
Even though there are many studies on spotting phishing using HTML, there are only a few recent datasets that anyone can use. Without Aljofey’s work [aljofey2022effective], it’s hard to find new HTML content datasets to compare how we’re doing. In our research, we collect and create data relevant to real-life phishing conditions. For detailed information about the dataset, please go to the Dataset section in our research (3.1). Additionally, we will share this dataset to contribute to the literature.
@@ 2 Related Work
Jain and Gupta (2016) introduced a client-side approach with an auto-updated white list, enhancing protection against phishing attacks. Their method provides individual users with an auto-updated white list, achieving an 86.02% true positive rate with less than 1.48% false negatives [jain2016novel]. Prakash et al. (2010) presented PhishNet, utilizing predictive blacklisting for phishing attack detection. PhishNet employs heuristics to identify simple modifications of known phishing sites, resulting in the discovery of around 18,000 new phishing URLs with low false positive (3%) and false negative (5%) rates [prakash2010phishnet]. Rao et al. (2020) developed CatchPhish, a lightweight application predicting URL legitimacy without website visits. The approach predicts URL legitimacy using features such as hostname, full URL, Term Frequency-Inverse Document Frequency (TF-IDF), and phish-hinted words, achieving high accuracy on benchmark datasets [rao2020catchphish]. Prabakaran et al. (2023) proposed an enhanced mechanism using variational autoencoders for identifying malicious URLs.
@@
Their approach integrates Variational Autoencoders (VAE) and deep neural networks (DNN) to automatically extract inherent features of raw URLs, achieving a maximum accuracy of 97.45% [prabakaran2023enhanced].
@@
URL-HTML Hybrid Approaches: Jain and Gupta (2019) proposed a machine learning-based approach for phishing detection, analyzing hyperlinks within the HTML source code. The approach introduces 12 categories of hyperlink-specific features, achieving an accuracy surpassing 98.4% on a logistic regression classifier [jain2019machine]. Rao et al. (2022) applied word embedding and machine learning to detect phishing websites, focusing on domain-specific text. Their multimodal approach achieved a remarkable accuracy of 99.34% using domain-specific text [rao2022wordembedding]. Aljabri et al. (2022) assessed lexical, network, and content-based features for detecting malicious URLs. The study identifies Naïve Bayes (NB) as the most effective model, achieving an accuracy of 96% [aljabri2022assessment]. Opara et al. (2020) introduced HTMLPhish, a deep learning-based approach for phishing web page detection through HTML analysis. HTMLPhish, using Convolutional Neural Networks (CNNs), achieved a high accuracy and True Positive Rate of over 93% [opara2020htmlphish]. Aljofey et al. (2022) presented a practical detection approach, combining URL and HTML features for phishing websites. Using character-level Term Frequency-Inverse Document Frequency (TF-IDF) features, their approach achieved high accuracy on both the authors’ dataset and a benchmark dataset [aljofey2022effective].
@@
Pretrained-based Approaches: Haynes et al. (2021) addressed phishing detection on mobile devices using NLP transformers, demonstrating lightweight URL-based phishing detection. Leveraging BERT and ELECTRA transformers, their approach showcases advantages in minimal training time, ease of updates, and real-time applicability on mobile devices [haynes2021lightweight]. Wang et al. (2023) proposed a lightweight multi-view learning approach using transformers with a mixture of experts for effective phishing attack detection. The method incorporates URL analysis, attributes, content, and behavioral information, outperforming state-of-the-art approaches [wang2023lightweight]. Misra and Rayz (2022) adapted pretrained language models to detect phishing emails, showcasing improvements over prior work. Their language models achieved nearly perfect performance on in-domain data and relative improvements on out-of-domain emails [misra2022lms].
@@
MLP-based Approaches: Odeh et al. (2020) developed an efficient prediction model using multilayer perceptron (MLP) to detect phishing websites. Their model aimed to efficiently predict phishing websites, contributing to improved internet security [odeh2020efficient]. Belfedhal and Belfedhal (2022) presented a lightweight phishing detection system based on machine learning and URL features, achieving high accuracy using MLP. The multilayer perceptron (MLP) model emerged as the most effective, showcasing robust performance in automatically detecting phishing web pages [belfedhal2022lightweight].
@@
Combined MLP+NLP Approaches: In recent years, the combination of Multilayer Perceptron (MLP) and Natural Language Processing (NLP) has demonstrated its effectiveness across various domains. This synergy has been particularly evident in the works of Yin et al. (2020) and Ramos et al. (2022). Yin et al. introduced TaBERT, a pretrained language model designed for semantic parsing over both textual and tabular data, showcasing the versatility of MLP and NLP integration [yin2020tabert]. Meanwhile, Ramos et al. leveraged BERT and numerical variables to classify injury leave based on accident descriptions, highlighting the potential of combining advanced NLP techniques with numerical features in a different domain [ramos2022combining].
@@
However, despite the success of MLP and NLP integration in various applications, the utilization of this combined approach in cybersecurity is very limited. Düzgün et al. (2023) proposed a network intrusion detection system that uniquely combines tabular and text-based features. Integrating deep learning and pretrained transformer methodologies, their approach stands out for achieving superior accuracy in detecting network anomalies [duzgun2023network].
@@ 3.1 Dataset
Furthermore, we incorporated the CatchPhish dataset, a widely used resource in existing literature. This dataset serves as a basis for comparing approaches, particularly in Aljofey’s work [aljofey2022effective], allowing us to assess the performance of our approach to others, with a specific emphasis on the analysis of HTML content. Table 1 provides a detailed overview of the data distribution within our research and CatchPhish datasets.
Table 1: Data Distribution in Research and CatchPhish Datasets
@@
Research Dataset 36,848 28,747 65,595
CatchPhish Dataset 32,972 27,280 60,252
3.2 Dataset Access
The research dataset is available for access on Google Drive as MTLP Dataset [dataset_reference].
@@ 3.3 Feature Extraction
Our webpage analysis employs an extensive feature extraction process, encompassing key information from various aspects. In the realm of textual features (C1, C2), ”page_title” and ”page_content” offer insights into the webpage’s context and main content, crucial for Natural Language Processing (NLP) models (NLP-1 and NLP-2).
On the numeric side, features from C3 to C13 introduce diverse characteristics. For example, ”hyperlink_count” (C3) quantifies the total number of hyperlinks, ”internal_count” (C4) specifically tallies internal links, and ”external_count” (C5) enumerates external links. Additionally, C6 to C9 encapsulates counts of external CSS and JavaScript files, identification of foreign external CSS, and counts of obfuscated JavaScript files, contributing collectively to a unified numeric analysis.
The feature ”suspicious_form_link” (C10) under login form features indicates links associated with a suspicious login form. Furthermore, C11 to C13 encompass aspects related to page structure and content, including ”common_page_ratio,” ”common_page_footer,” and ”has_meta_description.” These numeric features collectively provide a comprehensive view, aligning with a Multilayer Perceptron (MLP) model for integrated and effective webpage assessment.
The extraction of these features draws inspiration from several studies [aljofey2022effective] [opara2020htmlphish] [jain2019machine]. Table 2 illustrates the categories of features, their names, and the models to which they are related.
@@ 3.4.1 Multilayer Perceptron (MLP)
An MLP is an artificial neural network consisting of multiple layers of nodes [taud2018mlp], also known as neurons or perceptrons. It comprises an input layer, one or more hidden layers, and an output layer. Neurons in one layer are connected to neurons in the next layer by weights, and each connection may have an associated bias [aitkin2003statistical]. The activation function introduces non-linearity to the model, allowing it to learn complex patterns in the data.
@@ 3.4.2 Pretrained NLP Models
Transformer models, introduced by Vaswani et al. in 2017 [vaswani2017attention], have significantly advanced neural network architecture, particularly in Natural Language Processing (NLP). Their innovative use of self-attention mechanisms enables dynamic weighing of input elements, capturing long-range dependencies and relationships in sequential data like text or speech. These transformers are widely adopted across NLP applications, excelling in language translation, text classification (sentiment analysis, topic classification, document categorization), and named entity recognition.
Pretrained transformer models, a cornerstone in NLP, come pretrained on extensive datasets, offering general language representations. They serve as a foundation for fine-tuning specific tasks, providing benefits such as capturing intricate linguistic patterns as starting points for new models and facilitating efficient training by leveraging pretrained weights.
ALBERT (A Lite BERT) stands out as an efficient variant in the realm of pretrained transformer models [lan2019albert]. It addresses efficiency concerns without compromising performance through techniques like factorized embedding parameterization and cross-layer parameter sharing. BERT, a pioneering model, introduced bidirectional context to NLP, enhancing word relationships and semantics comprehension [devlin2018bert]. CANINE (Character Architecture with No Tokenization in Neural Encoders) is another pretrained encoder model designed to overcome the limitations of traditional tokenization techniques. Unlike conventional models, CANINE uses neural encoders that encode the sequence of characters or sub-words without explicitly tokenizing the input data [clark2021canine].
RoBERTa [tan2022roberta-lstm] (Robustly Optimized BERT Approach with Pre-training) refines BERT’s successes by modifying training strategies and eliminating the sentence prediction objective, achieving state-of-the-art performance across diverse NLP benchmarks. In our research, we utilize these pretrained models to compare their effectiveness in handling and extracting information from different textual features within the dataset, specifically the page_title and page_content columns.
@@ 3.4.3 MultiText-LP
Tailored specifically for handling numeric and categorical data, the MultiText-LP model incorporates an MLP to uncover intricate relationships and patterns within the dataset effectively. At the same time, it integrates two distinct NLP pretrained models for processing text information. NLP-1 is dedicated to features like ’page_title,’ while NLP-2 handles ’page_content,’ as illustrated in Figure 1.
@@
The synergy of these diverse models within the MultiText-LP architecture brings unparalleled versatility and performance, surpassing the capabilities of individual models. This combination allows the MultiText-LP model to handle numeric, categorical, and text data adeptly, making it a powerful tool for addressing complex classification challenges.
@@ 4.1 Evaluation Metrics
The F1-score, being the harmonic mean of precision and recall, ensures a balanced evaluation, critical in phishing detection where both minimizing false positives (precision) and capturing a significant portion of actual phishing instances (recall) are imperative. This metric is especially pertinent in the high-stakes nature of phishing identification, where the consequences of misclassifying benign content or overlooking genuine threats are substantial.
We also consider accuracy a secondary metric, given its importance in measuring overall performance. While accuracy provides a general assessment of correct predictions, the F1-score takes precedence in our evaluation, offering a more nuanced and targeted measure of our model’s efficacy in the specific challenge of phishing detection.
@@
Accuracy $\frac{TP+TN}{TP+TN+FP+FN}\times 100$
Percentage of correctly classified phishing and benign websites out of the total instances.
Precision $\frac{TP}{TP+FP}$
Proportion of correctly predicted phishing websites among those predicted as phishing.
Recall $\frac{TP}{TP+FN}$
Proportion of correctly predicted phishing websites out of all actual phishing websites.
[…]
Harmonic mean of precision and recall, providing a balanced measure that considers both false positives and false negatives.

### v3

@@ Abstract
The research findings highlight the effectiveness of the proposed approach, with the CANINE demonstrating superior performance in analyzing page titles and the RoBERTa excelling in evaluating page content. The fusion of two NLP and one MLP model, termed MultiText-LP, achieves impressive results, yielding a 96.80 F1 score and a 97.18 accuracy score on our research dataset. Furthermore, our approach outperforms existing methods on the Aljofey’s HTML dataset, showcasing its efficacies.
@@ 1 Introduction
The Internet Revolution began in the late 20th century and has significantly changed how we communicate, work, and live [1]. It has connected people worldwide, providing unprecedented access to information, resources, and opportunities. However, along with the benefits, there is a growing challenge of internet dependence and an increase in cyber-attacks [2].
One notable threat in this digital age is phishing, a deceitful practice where cybercriminals trick individuals into revealing sensitive information like usernames, passwords, or financial details [3]. Phishing techniques have become sophisticated, utilizing deceptive emails, fake websites, and malicious software. Staying vigilant and informed is crucial as cybercriminals continuously evolve their strategies.
Phishing has become the most common form of cybercrime today. According to the Anti-Phishing Working Group (APWG), in the second quarter of 2023, 1,286,208 phishing attacks were observed—the third-highest quarterly total on record [4]. The average wire transfer amount in Business Email Compromise (BEC) attacks surged to \$293,359, a 57% increase from Q1. Additionally, an estimated 3.4 billion phishing emails are sent daily, with 1,350,037 reported attacks in Q4 2022 alone. Notably, 95% of social engineering attack motivations are financially driven, emphasizing the high stakes involved [4].
Various software-based detection techniques are utilized in the ongoing fight against phishing attacks. Blacklists [5] and whitelists [6] serve as standard tools, but their effectiveness may be limited in identifying new or subtly modified phishing sites. Machine learning-based approaches leverage heuristic features like URL [7] [8], HTML content [9] [10] [11], website traffic, search engine data, and WHOIS records for classification, enhancing overall efficiency. Third-party services contribute valuable data to the detection process, including page rank, search engine indexing, and WHOIS [12] [13]. Visual similarity-based heuristics compare new websites with pre-stored signatures based on screenshots, font styles, images, page layouts, and logos [14] [15].
Even though there are many studies on spotting phishing using HTML, there are only a few recent datasets that anyone can use. Without Aljofey’s work [16], it’s hard to find new HTML content datasets to compare how we’re doing. In our research, we collect and create data relevant to real-life phishing conditions. For detailed information about the dataset, please go to the Dataset section in our study (3.1). Additionally, we will share this dataset to contribute to the literature.
@@ 2 Related Work
Jain and Gupta (2016) introduced a client-side approach with an auto-updated white list, enhancing protection against phishing attacks. Their method provides individual users with an auto-updated white list, achieving an 86.02% true positive rate with less than 1.48% false negatives [6]. Prakash et al. (2010) presented PhishNet, utilizing predictive blacklisting for phishing attack detection. PhishNet employs heuristics to identify simple modifications of known phishing sites, resulting in the discovery of around 18,000 new phishing URLs with low false positive (3%) and false negative (5%) rates [5]. Rao et al. (2020) developed CatchPhish, a lightweight application predicting URL legitimacy without website visits. The approach predicts URL legitimacy using features such as hostname, full URL, Term Frequency-Inverse Document Frequency (TF-IDF), and phish-hinted words, achieving high accuracy on benchmark datasets [7]. Prabakaran et al. (2023) proposed an enhanced mechanism using variational autoencoders for identifying malicious URLs.
@@
Their approach integrates Variational Autoencoders (VAE) and deep neural networks (DNN) to automatically extract inherent features of raw URLs, achieving a maximum accuracy of 97.45% [8].
@@
URL-HTML Hybrid Approaches: Jain and Gupta (2019) proposed a machine learning-based approach for phishing detection, analyzing hyperlinks within the HTML source code. The approach introduces 12 categories of hyperlink-specific features, achieving an accuracy surpassing 98.4% on a logistic regression classifier [9]. Rao et al. (2022) applied word embedding and machine learning to detect phishing websites, focusing on domain-specific text. Their multimodal approach achieved a remarkable accuracy of 99.34% using domain-specific text [10]. Aljabri et al. (2022) assessed lexical, network, and content-based features for detecting malicious URLs. The study identifies Naïve Bayes (NB) as the most effective model, achieving an accuracy of 96% [17]. Opara et al. (2020) introduced HTMLPhish, a deep learning-based approach for phishing web page detection through HTML analysis. HTMLPhish, using Convolutional Neural Networks (CNNs), achieved high accuracy and a true positive rate of over 93% [11]. Aljofey et al. (2022) presented a practical detection approach, combining URL and HTML features for phishing websites. Using character-level Term Frequency-Inverse Document Frequency (TF-IDF) features, their approach achieved high accuracy on both the authors’ dataset and a benchmark dataset [16].
@@
Pretrained-based Approaches: Haynes et al. (2021) addressed phishing detection on mobile devices using NLP transformers, demonstrating lightweight URL-based phishing detection. Leveraging BERT and ELECTRA transformers, their approach showcases advantages in minimal training time, ease of updates, and real-time applicability on mobile devices [18]. Wang et al. (2023) proposed a lightweight multi-view learning approach using transformers with a mixture of experts for effective phishing attack detection. The method incorporates URL analysis, attributes, content, and behavioral information, outperforming state-of-the-art approaches [19]. Misra and Rayz (2022) adapted pretrained language models to detect phishing emails, showcasing improvements over prior work. Their language models achieved nearly perfect performance on in-domain data and relative improvements on out-of-domain emails [20].
@@
MLP-based Approaches: Odeh et al. (2020) developed an efficient prediction model using multilayer perceptron (MLP) to detect phishing websites. Their model aimed to efficiently predict phishing websites, contributing to improved internet security [21]. Belfedhal and Belfedhal (2022) presented a lightweight phishing detection system based on machine learning and URL features, achieving high accuracy using MLP. The multilayer perceptron (MLP) model emerged as the most effective, showcasing robust performance in automatically detecting phishing web pages [22].
@@
Combined MLP+NLP Approaches: In recent years, the combination of Multilayer Perceptron (MLP) and Natural Language Processing (NLP) has demonstrated its effectiveness across various domains. This synergy has been particularly evident in the works of Yin et al. (2020) and Ramos et al. (2022). Yin et al. introduced TaBERT, a pretrained language model designed for semantic parsing over both textual and tabular data, showcasing the versatility of MLP and NLP integration [23]. Meanwhile, Ramos et al. leveraged BERT and numerical variables to classify injury leave based on accident descriptions, highlighting the potential of combining advanced NLP techniques with numerical features in a different domain [24].
@@
However, despite the success of MLP and NLP integration in various applications, the utilization of this combined approach in cybersecurity is very limited. Düzgün et al. (2023) proposed a network intrusion detection system that uniquely combines tabular and text-based features. Integrating deep learning and pretrained transformer methodologies, their approach stands out for achieving superior accuracy in detecting network anomalies [25].
@@ 3.1 Dataset
Furthermore, we incorporated the Aljofey’s dataset, a widely used resource in existing literature. This dataset serves as a basis for comparing approaches, particularly in Aljofey’s work [16], allowing us to assess the performance of our approach to others, with a specific emphasis on the analysis of HTML content. Table 1 provides a detailed overview of the data distribution within our research and Aljofey’s dataset. The research dataset is available on Google Drive as MTLP Dataset [26].
Table 1: Data Distribution in the MTLP and Aljofey’s Datasets
@@
MTLP Dataset [26] 36,848 28,747 65,595
Aljofey’s Dataset [16] 32,972 27,280 60,252
@@ 3.2 Feature Extraction
Our webpage analysis employs an extensive feature extraction process, encompassing key information from various aspects. In the realm of textual features (C1, C2), “page_title” and “page_content” offer insights into the webpage’s title and content, crucial for Natural Language Processing (NLP) models (NLP-1 and NLP-2).
On the numeric side, features from C3 to C13 introduce diverse characteristics. For example, “hyperlink_count” (C3) quantifies the total number of hyperlinks, “internal_count” (C4) specifically tallies internal links, and “external_count” (C5) enumerates external links. Additionally, C6 to C9 encapsulates counts of external CSS and JavaScript files, identification of foreign external CSS, and counts of obfuscated JavaScript files, contributing collectively to a unified numeric analysis.
The feature “suspicious_form_link” (C10) under login form features indicates links associated with a suspicious login form. Furthermore, C11 to C13 encompass aspects related to page structure and content, including “common_page_ratio,” “common_page_footer,” and “has_meta_description.” These numeric features collectively provide a comprehensive view, aligning with a Multilayer Perceptron (MLP) model for integrated and effective webpage assessment.
The extraction of these features draws inspiration from several studies [16] [11] [9]. Table 2 illustrates the categories of features, their names, and the models to which they are related.
@@ 3.3.1 Multilayer Perceptron (MLP)
An MLP is an artificial neural network consisting of multiple layers of nodes [27], also known as neurons or perceptrons. It comprises an input layer, one or more hidden layers, and an output layer. Neurons in one layer are connected to neurons in the next layer by weights, and each connection may have an associated bias [28]. The activation function introduces non-linearity to the model, allowing it to learn complex patterns in the data.
@@ 3.3.2 Pretrained NLP Models
Transformer models, introduced by Vaswani et al. in 2017 [29], have significantly advanced neural network architecture, particularly in Natural Language Processing (NLP). Their innovative use of self-attention mechanisms enables dynamic weighing of input elements, capturing long-range dependencies and relationships in sequential data like text or speech. These transformers are widely adopted across NLP applications, excelling in language translation, text classification (sentiment analysis, topic classification, document categorization), and named entity recognition.
Pretrained transformer models, a cornerstone in NLP, come pretrained on extensive datasets, offering general language representations. They serve as a foundation for fine-tuning specific tasks. They provide benefits such as capturing intricate linguistic patterns as starting points for new models and facilitating efficient training by leveraging pretrained weights.
ALBERT (A Lite BERT) stands out as an efficient variant in the realm of pretrained transformer models [30]. It addresses efficiency concerns without compromising performance through techniques like factorized embedding parameterization and cross-layer parameter sharing. BERT, a pioneering model, introduced bidirectional context to NLP, enhancing word relationships and semantics comprehension [31]. CANINE (Character Architecture with No Tokenization in Neural Encoders) is another pretrained encoder model designed to overcome the limitations of traditional tokenization techniques. Unlike conventional models, CANINE uses neural encoders that encode the sequence of characters or sub-words without explicitly tokenizing the input data [32].
RoBERTa [33] (Robustly Optimized BERT Approach with Pre-training) refines BERT’s successes by modifying training strategies and eliminating the sentence prediction objective, achieving state-of-the-art performance across diverse NLP benchmarks. In our research, we utilize these pretrained models to compare their effectiveness in handling and extracting information from different textual features within the dataset, specifically the page_title and page_content columns.
@@ 3.3.3 MultiText-LP
Tailored specifically for handling numeric and categorical data, the MultiText-LP model incorporates an MLP to effectively uncover intricate relationships and patterns within the dataset. At the same time, it integrates two distinct NLP pretrained models for processing text information. NLP-1 is dedicated to features like ‘page_title,’ while NLP-2 handles’ page_content,’ as illustrated in Figure 1.
@@
The synergy of these diverse models within the MultiText-LP architecture brings unparalleled versatility and performance, surpassing the capabilities of individual models. This combination allows the MultiText-LP model to adeptly handle numeric, categorical, and text data, making it a powerful tool for addressing complex classification challenges.
@@ 4.1 Evaluation Metrics
The F1-score, the harmonic mean of precision and recall, ensures a balanced evaluation. This is critical in phishing detection, where minimizing false positives (precision) and capturing a significant portion of actual phishing instances (recall) is imperative. This metric is especially pertinent in the high-stakes nature of phishing identification, where the consequences of misclassifying benign content or overlooking genuine threats are substantial.
Given its importance in measuring overall performance, we consider accuracy a secondary metric. While accuracy provides a general assessment of correct predictions, the F1-score takes precedence in our evaluation, offering a more nuanced and targeted measure of our model’s efficacy in the specific challenge of phishing detection.
@@
Accuracy $\frac{TP+TN}{TP+TN+FP+FN}\times 100$ Percentage of correctly classified phishing and benign websites out of the total instances.
Precision $\frac{TP}{TP+FP}$ Proportion of correctly predicted phishing websites among those predicted as phishing.
Recall $\frac{TP}{TP+FN}$ Proportion of correctly predicted phishing websites out of all actual phishing websites.
F1-Score $2\times\frac{\textit{Precision}\times\textit{Recall}}{\textit{Precision}+\textit{Recall}}$ Harmonic mean of precision and recall, providing a balanced measure that considers both false positives and false negatives.